OpenStudy (mollykenney):

sketch the graph and plot at least 5 points f(x) = -x^2 + 4x - 3

OpenStudy (immanuelv):

do you have a graphing calculator?

OpenStudy (mollykenney):

no :/

OpenStudy (anonymous):

i have a graphing calculator:)

OpenStudy (immanuelv):

*shocked* well what we can do is plot 5 points f(-2) = -(-2)^2 +4(-1) -3 f(-1) = -(-1)^2 +4(-1) -3 f(0) = -(0)^2 +4(0) -3 f(1) = -(1)^2 +4(1) -3 f(2) = -(2)^2 +4(2) -3

OpenStudy (anonymous):

nicceeee

OpenStudy (immanuelv):

awesome @Candlelover

OpenStudy (immanuelv):

@mollykenney does that help a bit mate?

OpenStudy (mollykenney):

i'm just confused with the plotting like i can draw the parabola but what are they looking for with the 5 points like… -3,0… 1,0… 3,0…2,1??

OpenStudy (anonymous):

what kind of math are u guys doing?

OpenStudy (immanuelv):

this is algebra

OpenStudy (mollykenney):

yes it does @immanuelv thank you so much

OpenStudy (anonymous):

Algerbra 1

OpenStudy (mollykenney):

2

OpenStudy (immanuelv):

@mollykenney glad that helped :D

OpenStudy (anonymous):

ohh

OpenStudy (mollykenney):

so is that what there looking for with the 5 points? like (-3,0) (1,0) ?

OpenStudy (immanuelv):

yeah

OpenStudy (anonymous):

ur very smart uv earned 89 meadls in math dang, oh well

OpenStudy (mollykenney):

ugh i feel so stupid that i'm confused… do you pick you own 5 numbers and plug them in when you first start the equation ? example how you chose -2,-1, etc

OpenStudy (immanuelv):

you got it @mollykenney you pick 5 numbers (since it wasn't given) and then you plug in to x to find f(x) aka y

OpenStudy (mollykenney):

wow thank you so much @immanuelv i really appreciate it

OpenStudy (immanuelv):

@mollykenney your're welcome :D